7^{th} International Conference on the Discrete Simulation of Fluids University of Oxford, 14th-18th July 1998

Registration Form

Full Name: Affiliation: Address: E-mail: Telephone: Fax: Special dietary or other requirements:

Accommodation (please tick as appropriate)

Monday 13th July  □ bed & breakfast en-suite @ £43

□ bed & breakfast standard @ £30

Tuesday 14th July  □ dinner @ £15

□ bed & breakfast en-suite @ £43 □ bed & breakfast standard @ £30

Wednesday 15th July □ dinner @ £15

□ bed & breakfast en-suite @ £43 □ bed & breakfast standard @ £30

Thursday 16th July conference dinner (included in registration)

□ bed & breakfast en-suite @ £43 □ bed & breakfast standard @ £30

Friday 17th July  □ dinner @ £15

□ bed & breakfast en-suite @ £43 □ bed & breakfast standard @ £30

Saturday 18th July  □ bed & breakfast en-suite @ £43

□ bed & breakfast standard @ £30

Registration (please tick as appropriate)

Standard  □ £150 before March 15th

□ £250 thereafter

Student  □ £100 before March 15th

□ £200 thereafter

Students should ask their supervisor to sign their application form

TOTAL  £

Please send with a cheque in Pounds Sterling or notification of bank transfer to:

Dr. Julia Yeomans Department of Theoretical Physics 1, Keble Road Oxford OX1 3NP UK